STARK SPECTROSCOPY OF CH_{3}F SOLVATED IN HELIUM NANODROPLETS

\underline{ALEXANDER M. MORRISON}, STEVEN D. FLYNN, TAO LIANG and GARY E. DOUBERLY, Department of Chemistry, University of Georgia, Athens, Georgia 30602.

Infrared spectra in the 2800-3150 cm^{−1} region are presented for CH_{3}F embedded in helium nanodroplets. We observe both the ν_{1}(a_{1}) symmetric C-H stretch and the ν_{4}(e) asymmetric C-H stretch. Transitions from K =0(A) and K =1(E) are observed due to nuclear spin statistics. The symmetric stretch is split due to a Fermi resonance between ν_{1} and 2ν_{5} (overtone of the asymmetric bend). The (ν_{4},2ν_{5}) Fermi resonance is an A_{1}-A_{1} type parallel band and ν_{4} is a E-A_{1} type perpendicular transition. The upper diad of the (ν_{1},2ν_{5}) Fermi resonance, found around 2960 cm^{−1} is homogeneously broadened due to efficient vibrational relaxation to the lower diad. Furthermore, the J =2, K =2 ← J =1, K =1 transition (^{r}R_{1}(1)) is significantly broadened due to rotational resonances with the phonon and roton modes of the helium droplet. Further attempts to observe broadening effects are studied by measuring the Stark Spectra of all three bands at many electric field strengths.